Grade 10Electric Current

electric powar

prateek  nagar
14 Years agoGrade 10

2 Answers

Hasan  Ali
14 Years ago

Electeric power is ratio of electric work by time

 

         P = W/t

 

 

 

Plz....... Cilick on YES.......

Ashwin Sinha
14 Years ago

Dear Prateek Nagar,

Electric power is the rate at which electric energy is transferred by an electric circuit. The SI unit of power is the watt.

When electric current flows in a circuit, it can transfer energy to do mechanical orthermodynamic work. Devices convert electrical energy into many useful forms, such as heat(electric heaters), light (light bulbs), motion (electric motors), sound (loudspeaker), information technological processes (computers), or even chemical changes. Electricity can be produced mechanically by generation, or chemically, or by direct conversion from light inphotovoltaic cells, also it can be stored chemically in batteries.

In direct current resistive circuits, electrical power is calculated using Joule's law:

P = IV \,

where P is the electric power, V the potential difference, and I the electric current.

In the case of resistive (Ohmic, or linear) loads, Joule's law can be combined with Ohm's law (I = V/R) to produce alternative expressions for the dissipated power:

P = I^2 R = \frac{V^2}{R},

where R is the electrical resistance.

 

 

All The Best............

 

 

Hope this helped you dear........

 

Plz. approve my answer by clicking 'Yes' given below, if you loved it......... Plz. Plz. Plz. don't forget to do it.......Smile